OpenStudy (henrietepurina):

Greg has only nickels and dimes in his money box. He knows that he has more than $10 in the box. Let x represent the number of nickels in the box and y represent the number of dimes in the box. Which of the following statements best describes the steps to graph the solution to the inequality in x and y?

OpenStudy (henrietepurina):

Draw a dashed line to represent the graph of 5x + 10y = 1000, and shade the portion below the line for positive values of x and y. Draw a dashed line to represent the graph of 5x + 10y = 1000, and shade the portion above the line for positive values of x and y. Draw a dashed line to represent the graph of 10x + 5y = 1000, and shade the portion above the line for positive values of x and y. Draw a dashed line to represent the graph of 10x + 5y = 1000, and shade the portion below the line for positive values of x and y.

undeadknight26 (undeadknight26):

I would say B or C since they are the same answer choices...since they know that they have MORE than $10 in the jar it would have to be shaded above the line mate!

undeadknight26 (undeadknight26):

Study on!

OpenStudy (henrietepurina):

Thanks :)

undeadknight26 (undeadknight26):

Any time mate!

OpenStudy (henrietepurina):

How can I find out B or C?

undeadknight26 (undeadknight26):

well...If you check each question..they have the same EXACT words in them...and say the EXACT same thing...so honestly they look like they posted 2 of the same question answers...

OpenStudy (henrietepurina):

Dang, what can I do....

undeadknight26 (undeadknight26):

I would say take a small screen shot of this question or snipping tool if you have windows 7...and if u get it wrong...ask a teacher about it and show her the evidence you have :D

OpenStudy (henrietepurina):

What if i said C?

OpenStudy (henrietepurina):

Yeah I like your idea....

undeadknight26 (undeadknight26):

Also when taking the screen shot...I recommend you take this or any other tabs you have open such as yahoo answers etc. You could get in trouble for using these sites for help for school :D

OpenStudy (henrietepurina):

yeah i know, thanks a lot :)

undeadknight26 (undeadknight26):

Any time mate!

OpenStudy (henrietepurina):

can you help on a few more?

undeadknight26 (undeadknight26):

Sure!

OpenStudy (henrietepurina):

What is the solution to the equation?

OpenStudy (henrietepurina):

x = –2 x = –1 x = 1 x = 2

undeadknight26 (undeadknight26):

Not very good in these sorry mate...

OpenStudy (henrietepurina):

i say its 2

undeadknight26 (undeadknight26):

hmm try wolframalpha.com...